Value of (a–c)[(a–b) 2 +(b–c) 2 –(a–b)(b–c)]+(c–a) 3 is equal to (1) 3(a – b) (b – c) (c –a) (2) (a – b) (b – c) (c – a) (3)(a–b) 3 (b–c) 3 (c–a) 3 (4)None of these

  1. Value of (a–c)[(a–b)2 +(b–c)2 –(a–b)(b–c)]+(c–a)3 is equal to

    1. (1) 3(a – b) (b – c) (c –a) (2) (a – b) (b – c) (c – a)

      (3)(a–b)3 (b–c)3 (c–a)3 (4)None of these 

Grade:10

1 Answers

Arun
6 years ago

Solution

={(a-b)+(b-c) }{(a - b)^2 + (b - c)^2 - (a - b) (b - c)} + (c - a)^3

(x+y) (x^2 - xy+y^2) =x^3 + y^3

=(a-b)^3 +(b-c)^3 +(a-c) ^3

On solving it further

=0